An ice-making machine operates in a Carnot cycle. It takes heat from water at 0.0 ∘C and rejects heat to a room of temperature 26.0 ∘C. Suppose that a mass 82.0 kg of water at 0.0 ∘C is converted to ice at 0.0 ∘C.

a)How much heat is rejected to the room?

b)How much energy must be supplied to the device?
Answer #1

Here ,

Tc = 0 degree C = 273 K

Th = 26 degree C = 299 K

heat taken from cold reserviour , Qc = 82 * 334 KJ

Qc = 27388 KJ

a) for the heat rejected

Qh/Qc = Th/Tc

Qh/27388 = 299/273

Qh = 29996.4 KJ

the heat rejected to the room is 29996.4 KJ

b)

energy supplied to device = Qh - Qc

energy supplied to device = 29996.4 - 27388

energy supplied to device = 2608.4 KJ

the energy supplied to device is 2608.4 KJ